# Line Search Strategy for Navigating through Barren Plateaus in Quantum Circuit Training

### v2

@@ Discussion
In this manuscript, we have presented a novel optimization method for training quantum circuits. A prevalent challenge in variational quantum algorithms is the hindering impact of barren plateaus (BPs) on circuit training. The vanishing gradient problem associated with BPs significantly limits the applicability of near-term quantum algorithms. Clearly speaking, the application of a given variational algorithm becomes questionable when the training process fails at large scaled problem instances. Therefore, developing optimization strategies resistant to BPs is highly desirable. We demonstrated that the long-range properties of the optimization landscape can be leveraged to determine a search direction and optimization bounds along this direction. This method introduces a novel line-search approach for training quantum circuits, autonomously maintaining low values of the second-Rényi entropy during optimization. Given the inherent connection between entropy and BPs, this property of the developed algorithm implies a natural tendency to avoid BPs during training without requiring external control mechanisms.
We applied the optimization strategy to VQE simulations on randomly generated Heisenberg XXX and Sachdev–Ye–Kitaev (SYC) models, all of which were previously identified to exhibit barren plateaus (BPs) during VQE process. We increased the depth of the training circuits to $p=1000$ entangling layers with $16$ qubits. The developed optimization strategy, utilizing the gradient-free search direction method, consistently decreased the VQE energy over thousands of iterations without getting trapped by BPs. We demonstrated these outstanding capabilities of the solver by training quantum circuits containing up to $15000$ entangling gates.

### v3

@@ Discussion
In this manuscript, we have presented a novel optimization method for training quantum circuits. A prevalent challenge in variational quantum algorithms is the hindering impact of BPs on circuit training. The vanishing gradient problem associated with BPs significantly limits the applicability of near-term quantum algorithms. Clearly speaking, the application of a given variational algorithm becomes questionable when the training process fails at large scaled problem instances. Therefore, developing optimization strategies resistant to BPs is highly desirable. We demonstrated that the long-range properties of the optimization landscape can be leveraged to determine a search direction and optimization bounds along this direction. This method introduces a novel batched line-search approach for training quantum circuits, autonomously maintaining low values of the second-Rényi entropy during optimization, provided the optimization starts with zero qubit rotation initialization. Given the inherent connection between entropy and BPs, this property of the developed algorithm implies a natural tendency to avoid BPs during training without requiring external control mechanisms. This is achieved by effectively reducing the expressibility (or variability) of the circuit ansatz via updating only a smaller subset of the optimization parameters in each iteration. Compared to traditional gradient based optimization approaches – like the plain gradient descent or the ADAM algorithm – our numerical results yielded an increased efficiency of our optimization approach during the first periods of the optimization process, where our method showcased significantly larger improvements on the cost function. This is where our optimization method delivers advantage, as apparently bigger improvement can be achieved within fewer cost function evaluations. Powell’s gradient free optimization method is the one we should also mention here, as our approach shares some similarities with this method. Therefore, the numerical experiments performed with Powell’s optimization method got quite close to our results. Our approach is capable of outperforming Powell’s method by two means. Firstly, our method adopts the properties of the optimization landscape in making the updates of the parameters. Secondly, in contrast to Powell’s method, our approach can be parallelized, significantly shortening the training time.
We applied the optimization strategy to VQE simulations on randomly generated Heisenberg XXX and Sachdev–Ye–Kitaev (SYC) models, all of which were previously identified to exhibit barren plateaus (BPs) during the VQE process. We increased the depth of the training circuits to $p=1000$ entangling layers with $16$ qubits. The developed optimization strategy, utilizing the gradient-free search direction method, consistently decreased the VQE energy over thousands of iterations without getting trapped by BPs. We demonstrated these outstanding solver capabilities by training quantum circuits containing up to $15000$ entangling gates.
@@ Acknowledgments
Appendix A Technical details on the derivation of Eqs. (5) and (7)
The derivation of Eqs. (5) and (7) relies on the analytical properties of the quantum gates incorporated in the trained circuits. In general, each of the parametric gates can be expressed via trigonometric functions of its parameter, namely:
$U=\cos(p)U_{1}+\sin(p)U_{2}+U_{3}\;,$ (19)
where constant matrices $U_{i}$ are of the same size as the matrix of the gate operation labeled by $U$. For single qubit rotations (including RX, RY, RZ) the matrix $C$ is zero, while for constant gates (like the $CNOT$ gate) matrices $A$ and $B$ are zeros. By applying this gate on an initial state $|0\rangle$, we arrive at the transformed state
$|\Psi\rangle=U_{4}\times\bigg{(}\cos(p)U_{1}+\sin(p)U_{2}+U_{3}\bigg{)}\times U_{5}|0\rangle\;,$ (20)
where $U_{i}$’s are constant matrices with respect to the chosen parameter $p$. This expression can be further simplified to
$|\Psi\rangle=\cos(p)|a\rangle+\sin(p)|b\rangle+|c\rangle$ (21)
with respect to the parameter $p$, where $|a\rangle$, $|b\rangle$ and $|c\rangle$ are suitable states from the Hilbert space. By evaluating the expectation value of the Hamiltonian we get:
$\langle\Psi|\hat{H}|\Psi\rangle$ $=$ $\cos(p)^{2}H_{aa}+\sin(p)^{2}H_{bb}+H_{cc}+$ (22)
$+$ $\cos(p)\sin(p)\left(H_{ab}+H_{ba}\right)+$ (23)
$+$ $\cos(p)(H_{ac}+H_{ca})+\sin(p)(H_{bc}+H_{cb})\;,$ (24)
where matrix elements $H_{ij}$ ($i,j\in\{a,b,c\}$) are defined as $H_{ij}=\langle i|\hat{H}|j\rangle$. After utilizing trigonometric identities $2\cos(p)^{2}=\cos(2p)+1$ and $2\sin(p)\cos(p)=\sin(2p)$, the equation simplifies to:
$\langle\Psi|\hat{H}|\Psi\rangle$ $=$ $\cos(2p)\frac{H_{aa}-H_{bb}}{2}+\sin(2p)\frac{H_{ab}+H_{ba}}{2}$ (25)
$+$ $\cos(p)H_{ac}+\sin(p)H_{bc}+\frac{H_{ba}+H_{bb}}{2}+H_{cc}\;.$ (26)
Since we can always transform the expression $\cos(p)c_{1}+\sin(p)c_{2}$ into a form of $A\cos(p+\varphi)$ with $A=\sqrt{(c_{1})^{2}+(c_{2})^{2}}$ and $\sin(\varphi)=c_{1}/A$, we can transform the equation above into the form of
$E_{VQE}=\kappa\cdot\sin(2p_{i}+\xi)+\gamma\cdot\sin(p_{i}+\phi)+C\;,$ (27)
were the parameters $\kappa$, $\xi$, $\gamma$, $\phi$ and $C$ are determined by the Hamiltonian elements $H_{ij}$ ($i,j\in\{a,b,c\}$). When the (19) representation of the selected gate contains no constant $U_{3}$, then we arrive at the expression:
$E_{VQE}=\kappa\cdot\sin(2p_{i}+\xi)+C\;.$ (28)
These equations match Eqs. (5) and (7) in the main text.